# Which chromium-sulfur compounds exist as 2D material?

### v1

@@ 3 Conclusion
We demonstrate the growth of two new chromium-sulfur 2D materials, Cr2S3-2D and CrS4-2D, in single unit cell thickness. Cr2S3-2D exhibits threefold symmetry with an in-plane lattice constant of nm, an apparent STM height of 0.79 nm, and 2 Cr atoms per unit cell. Upon annealing to 850 K or growing over the complete layer, Cr2S3-2D transforms into CrS4-2D with the same lattice parameter, but a thickness of 1.02 nm. This transition results in an additional -R30∘ superstructure in respect to the original lattice. Both materials can also be grown phase pure to almost a full layer. Cr2S3-2D and CrS4-2D display a NiAs-type structure terminated by S planes. However, the middle Cr layer in CrS4-2D displays Cr vacancies with 1/3 of the Cr positions missing, resulting in a ()-R30° superstructure.
@@
Our DFT calculations predict that both 2D materials are indirect bandgap semiconductors in their freestanding form. Interestingly, we find that when supported on Gr, they exhibit metallic behavior due to their relatively large work function compared to Gr. DFT calculations find sizable magnetic moments for both phases, possessing magnetic order with coupling depending on the degree of Coulomb repulsion, a characteristic that requires experimental validation. The phase-pure, large-area growth of Cr2S3-2D and CrS4-2D enables detailed investigation of their magnetic properties using averaging techniques and establishes these compounds as promising 2D magnetic materials for potential applications.
4 Experimental Section
4.0.1 Experimental methods
In situ growth and analysis of CrxSy-2D were performed in a variable temperature and a low temperature bath cryostat ultra-high vacuum STM system, both with base pressure of mbar. Each system is equipped with standard sample preparation tools, MBE growth facility, LEED or microchannel plate LEED, and STM operating down at 300 K and at 1.7 K.
4.0.2 Substrate preparation
@@
To grow a single crystal Gr on Ir(110), clean unreconstructed Ir(110) is heated to 1510 K and exposed at this temperature for 240 s to mbar ethylene 54. To grow single crystal Gr on Ir(111), cleaned Ir(111) is exposed to mbar ethylene at room temperature for 120 sec. After flash annealing to 1470 K without ethylene, subsequently the sample is exposed to mbar ethylene at 1370 K for 600 sec for completion of Gr growth 64. The quality of the as-grown Gr on Ir(110) and Ir(111) is checked by LEED and STM.
4.0.3 Sample preparation
To grow CrxSy-2D, Gr/Ir(110) is exposed to a Cr flux in the range of atoms m-2 s-1 from an e-beam evaporator simultaneously with a sulfur flux obtained by decomposing pyrite (FeS2) in a Knudsen cell in a distance of 8 cm from the sample. The S flux is characterized by a distant ion gauge measuring a S pressure of mbar. This growth step was performed typically at 300 K, if not specified differently. To complete CrxSy-2D formation, the sample is subsequently annealed for half the initial deposition time in the same sulfur flux at temperatures ranging from 500 K to 1050 K and specified in the respective captions.
4.0.4 LEED measurements
@@
4.0.5 STM measurements
STM data is acquired at 300 K in the variable temperature STM system and at 1.7 K in the bath cyrostat STM system. Constant current topographies are presented with sample bias and tunneling current specified in the captions. Image processing was conducted using WSxM 65.
4.0.6 STS measurements
@@
4.0.7 Theoretical calculations
The energetics and magnetic properties of all considered systems were investigated using spin-polarized density functional theory (DFT) as implemented in the VASP code 66, 67. All the calculations were carried out using the Perdew-Burke-Ernzerhof exchange-correlation functional 68. To account for the high correlation effects among d-orbital electrons of chromium atoms, an effective Hubbard value () was considered for Fe atoms within the DFT + U approach. A plane-wave cut-off energy of 500 eV and force tolerance of 0.01 eV/Å was set for geometry optimization. Van der Waals interactions were considered using the DFT-D2 method proposed by Grimme 69. A vacuum space of approximately 30 Å was introduced to model the slabs between the periodic images in the confinement direction.
@@ 4.0.7 Theoretical calculations
Funding from Deutsche Forschungsgemeinschaft (DFG) through CRC 1238 (project number 277146847, subprojects A01 and B06) is acknowledged. J.F. and W.J. acknowledge financial support from the DFG through project FI 2624/1-1 (project No. 462692705) within the SPP 2137 and project JO 1972/2-1 (project No. 535290457) within SPP 2244, respectively. A.V.K. acknowledges funding from DFG projects KR 4866/11-1, and the collaborative research center “Chemistry of Synthetic 2D Materials” CRC-1415-417590517. Generous CPU time grants from the Technical University of Dresden computing cluster (TAURUS) and Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu), Supercomputer HAWK at Höchstleistungsrechenzentrum Stuttgart (www.hlrs.de), are greatly appreciated.
Conflict of Interest

### v2

@@ III Conclusion
We demonstrate the growth of two new chromium-sulfur 2D materials, Cr2S3-2D and CrS4-2D, in single unit cell thickness. Cr2S3-2D exhibits threefold symmetry with an in-plane lattice constant of nm, an apparent STM height of 0.79 nm, and 2 Cr atoms per unit cell. Upon annealing to 850 K or growing over the complete layer, Cr2S3-2D transforms into CrS4-2D with the same lattice parameter, but a thickness of 1.02 nm. This transition results in an additional -R30∘ superstructure in respect to the original lattice. Both materials can also be grown phase-pure to almost a full layer. Cr2S3-2D and CrS4-2D display a NiAs-type structure terminated by S planes. However, the middle Cr layer in CrS4-2D displays Cr vacancies with 1/3 of the Cr positions missing, resulting in a ()-R30° superstructure.
A key factor to develop a proper structure model was the precise STM-based calibration of the amount of Cr per unit cell. This was accomplished by growing and measuring the area of Cr islands on Ir(111). While our method could not be applied to determine the S content in the new materials, we envision that the same type of calibration may be possible for the less volatile Se and Te. Thereby the full stoichiometry for 2D materials of unknown structure could be provided with precision, making our calibration method generally relevant in 2D material structure determination.
@@
Combining the STM and LEED data with DFT calculations enabled us to establish the structure of the two new 2D materials in detail. Remarkably, surface induced relaxation of the Cr layers is apparent and may be a factor stabilizing the new 2D materials. Our DFT calculations predict that both 2D materials are indirect bandgap semiconductors in their free-standing form. Interestingly, we find that when supported on Gr, they exhibit metallic behavior due to their relatively large work function compared to Gr. DFT calculations find sizable magnetic moments for both phases, possessing magnetic order with coupling depending on the degree of Coulomb repulsion, a characteristic that requires experimental validation. The phase-pure, large-area growth of Cr2S3-2D and CrS4-2D enables detailed investigation of their magnetic properties using averaging techniques and establishes these compounds as promising 2D magnetic materials for potential applications.
IV Experimental Section
IV.1 Experimental methods
In situ growth and analysis of CrxSy-2D were performed in a variable temperature and a low temperature bath cryostat ultra-high vacuum STM system, both with base pressure of mbar. Each system is equipped with standard sample preparation tools, MBE growth, LEED or microchannel plate LEED, and STM operating at 300 K and at 1.7 K.
IV.2 Substrate preparation
@@
To grow single crystal Gr on Ir(110), clean unreconstructed Ir(110) is heated to 1510 K and exposed at this temperature for 240 s to mbar ethylene [62]. To grow single crystal Gr on Ir(111), clean Ir(111) is exposed to mbar ethylene at room temperature for 120 sec. After flash annealing to 1470 K without ethylene, subsequently the sample is exposed to mbar ethylene at 1370 K for 600 sec for completion of Gr growth [76]. The quality of the as-grown Gr on Ir(110) and Ir(111) is checked by LEED and STM.
IV.3 Sample preparation
To grow CrxSy-2D, Gr/Ir(110) is exposed to a Cr flux in the range of atoms m-2 s-1 from an e-beam evaporator simultaneously with a sulfur flux obtained by decomposing pyrite (FeS2) in a Knudsen cell in a distance of 8 cm from the sample. The S flux is characterized by a distant ion gauge measuring a S pressure of mbar. This growth step was typically performed at 300 K, if not specified differently. To complete CrxSy-2D formation, the sample is subsequently annealed for half the initial deposition time in the same sulfur flux at temperatures ranging from 500 K to 1050 K and specified in the respective captions.
IV.4 LEED measurements
@@
IV.5 STM measurements
STM data is acquired at 300 K in the variable temperature STM system and at 1.7 K in the bath cyrostat STM system. Constant current topographies are presented with sample bias and tunneling current specified in the captions. Image processing was conducted using WSxM [77].
IV.6 STS measurements
@@
IV.7 Theoretical calculations
The energetics and magnetic properties of all considered systems were investigated using spin-polarized density functional theory (DFT) as implemented in the VASP code [78, 79]. All the calculations were carried out using the Perdew-Burke-Ernzerhof exchange-correlation functional [80]. To account for the high correlation effects among d-orbital electrons of chromium atoms, an effective Hubbard value () was considered for Cr atoms within the DFT + U approach [81]. A plane-wave cut-off energy of 500 eV and force tolerance of 0.01 eV/Å was set for geometry optimization. Van der Waals interactions were considered using the DFT-D2 method proposed by Grimme [82]. A vacuum space of approximately 30 Å was introduced to model the slabs between the periodic images in the confinement direction.
@@ IV.7 Theoretical calculations
Acknowledgements.
Funding from Deutsche Forschungsgemeinschaft (DFG) through CRC 1238 (project number 277146847, subprojects A01 and B06) is acknowledged. J.F. and W.J. acknowledge financial support from the DFG through project FI 2624/1-1 (project No. 462692705) within the SPP 2137 and project JO 1972/2-1 (project No. 535290457) within SPP 2244, respectively. A.V.K. acknowledges funding from DFG projects KR 4866/11-1, and the collaborative research center “Chemistry of Synthetic 2D Materials” CRC-1415-417590517. Generous CPU time grants from the Technical University of Dresden computing cluster (TAURUS) and Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu), Supercomputer HAWK at Höchstleistungsrechenzentrum Stuttgart (www.hlrs.de), are greatly appreciated.Conflict of Interest